# Optimizing wheel loader performance — an end-to-end approach

### v2

@@ Abstract
Wheel loaders in mines and construction sites repeatedly load soil from a pile to load receivers. This task presents a challenging optimization problem since each loading’s performance depends on the pile state, which depends on previous loadings. We investigate an end-to-end optimization approach considering future loading outcomes and transportation costs between the pile and load receivers. To predict the evolution of the pile state and the loading performance, we use world models that leverage deep neural networks trained on numerous simulated loading cycles. A look-ahead tree search optimizes the sequence of loading actions by evaluating the performance of thousands of action candidates, which expand into subsequent action candidates under the predicted pile states recursively. Test results demonstrate that, over a horizon of 15 sequential loadings, the look-ahead tree search is 6 % more efficient than a greedy strategy, which always selects the action that maximizes the current single loading performance, and 14 % more efficient than using a fixed loading controller optimized for the nominal case.
keywords:
journal: a journal
[…]
[AlgoryxAddress]organization=Algoryx Simulation,addressline=, city=Umeå, postcode=SE-90736, state=, country=Sweden
[…]
[KAUAddress]organization=Karlstad University,addressline=, city=Karlstad, postcode=SE-651 88, state=, country=Sweden
@@ 1 Introduction
The task of most construction and mining machines, such as wheel loaders, is to perform long sequences of earthmoving operations in dynamic environments. Optimizing this task involves selecting loading and transportation actions that maximize the gain over time, according to some performance measure. This is challenging for several reasons. First, the performance of a single loading cycle depends highly on the local shape of the pile and the soil properties [1]. The loading actions need to be carefully adapted to these conditions. Second, each loading alters the pile shape. From the perspective of future loadings, such a change can be either for the better or worse. There might be actions that are of low performance in the short term but are necessary to achieve higher gains in the future [2]. Therefore, a greedy strategy of always choosing the loading action that maximizes the performance for a single loading might be sub-optimal in the long run. Such a strategy might gradually deteriorate the pile state or lead to excessive transportation between the dig location and the load receiver. This motivates considering future states when selecting the actions.
We refer to the problem of finding the action sequence that maximizes the total performance of many wheel loading cycles, given the initial pile state, as the wheel loader end-to-end optimization problem. To the best of our knowledge, this problem has not been scientifically explored previously. In fact, until recently, it was computationally intractable. In our previous work [3], we developed data-driven models capable of predicting the loading performance and the subsequent state of the pile within a few milliseconds. Model inputs are the previous pile state and selected action parameters. This makes it possible to rapidly evaluate the net performance of numerous alternative action sequences in advance.
@@
We use a look-ahead tree search approach illustrated in Fig. 1. Future action candidates for the foreseen future states are identified recursively. The action candidates are evaluated, using data-driven models developed in previous research [3], to compute their respective performance and expand into further action candidates under the new states. This approach is, reportedly, similar to how experienced operators intuitively plan their work, but we expect that a computerized model and search algorithms can make more accurate predictions and plan over longer horizons. To understand the importance of the look-ahead search, we examined the effect of using different planning horizons and objectives.
@@
The wheel loader is assumed to have a high-level planner and a low-level control system, which work as the hierarchical architecture described in refs. [4] and [5]. The high-level planner selects a subtask action, such as bucket filling, dumping, or V-shaped transportation, according to the current state of the machine and the work site. A low-level controller performs the subtasks given a motion plan and control parameters. For bucket filling, the machine is equipped with a type of admittance controller [6]. Our controller takes four action parameters that determine how reactive the boom and bucket actuation is in response to the momentaneous dig force [3]. The performance of a loading cycle is measured in terms of mass, time, and work, with contributions from each subtask.
@@ 2 Related work
The end-to-end optimization problem of sequential loading was discussed in [2], considered intractable, and substituted with a simpler problem of optimizing each loading cycle by searching in a lower-dimensional action space implicitly constrained by the pile state. An average of 90 % efficiency (bucket payload relative to its capacity) was achieved over 20 cycles. The study was limited to 2D, used a heuristic simulation model for soil displacements and settling, and did not consider time or energy consumption. Attempts were made to improve the performance further by using heuristics to avoid action plans that produce challenging terrain configurations, but the optimization was computationally prohibitive.
@@
What is a good dig location in the pile, given its shape in terms of a heightmap, has been studied from the perspective of both bucket filling and the transportation between the dig and the dump location. In [7], a coarse-level planner finds the points along the pile contour closest to the load receiver, and a refined planner was used to avoid curved regions of the pile surface to minimize bucket-side loading and drop-in filling efficiency. A sequential coarse-to-fine planner was studied in [8] and compared to a greedy planner. The coarse planner first produces a sequence of dig regions ordered radially around the pile centroid. The fine planner then selects the dig points inside each region that have the best local pile shape. Simulations (cellular automata) of 50 consecutive loadings showed that the coarse-to-fine strategy maintains a good pile shape and performs at 80–90 % while a greedy strategy (always selecting the dig location with optimal pile shape) initially performs equal or better for ten loadings but drops down to about 60 % after 25 loadings.
@@
In [9], reinforcement learning was used to train a multiagent system for autonomous control of an underground load-haul-dump (LHD) vehicle using a depth camera, lidar, and force and kinematics sensors. One agent was used to select a favorable dig position, given the observed pile shape. A second agent had the task of steering the vehicle toward the target position and controlling the forward drive as well as the lift and tilt cylinders to perform bucket filling. The agents were trained in a simulated environment to learn control policies designed for high productivity and energy efficiency over multiple loading cycles while also avoiding collisions and wheel slip. The system achieved bucket filling with 75 % of maximum capacity on average and used 4 % less energy by actively selecting favorable dig locations. The vehicle was confined to operate in a narrow mine drift with little variability in viable dig position or how to navigate to and from the dump location. It is untested how well this approach applies to controlling a wheel loader above ground with much larger action and observation space.
Multiobjective optimization of an LHD control strategy was explored in [10] using surrogate models based on data from sequences of simulated loadings in a pile of fragmented rock. Each simulated sequence used a parameterized dig trajectory planner that adapts to the local pile state. Based on many simulations with different parameters, the learned model can identify the loading strategies that ensure high average performance and avoid those that systematically deteriorate the pile state.
@@
Optimization of the V-shaped transportation path between the dig and dump location has been addressed in several studies, focusing on finding the shortest path with curvature consistent with the vehicle’s minimum turning radius [11, 12], taking into account the vehicle dynamics and construction working site constraints [13], non-uniform trajectories under forward and reverse driving conditions [14], or minimizing fuel consumption [15]. Time and energy-efficient tracking of a planned path has been explored in several studies using optimal control [16, 17].
@@
The optimal bucket filling trajectory was researched in [18] using the discrete element method (DEM) and used in [19] to find the optimal short loading cycle using dynamic programming. For the sake of computational efficiency, the simulations were limited to quasi-2D piles sloped with the material’s angle of repose. The effect of the pile shape on the optimal trajectory or the resulting 3D pile state has not been explored using DEM simulation, apart from the work in [10].
@@ 4.1 Loading prediction
We use the data-driven world model developed in [3] for a Komatsu WA320-7 wheel loader doing automatic bucket filling in piles of gravel using an admittance controller. The models predict the loading outcome
@@
As in [3], we use different sizes and resolutions of the local heightmap for the pile state and performance predictor models. For the former model, we discretize a quadratic heightmap with 5.2 m sides in a $52\times 52$ grid. For the latter model, we use a $36\times 36$ grid and a side length of 3.6 m.
@@ 4.1.1 Optimal loading action parameters
where $\eta$ is the step length and $\nabla=\partial/\partial\bm{a}^{\text{load}}$. As an initial guess, we use $\bm{a}^{\text{load}}_{n}=[1,0,0.0,0.5,0.5]^{\mathsf{T}}$, which corresponds to the machine thrusting deep into the pile, filling the bucket with little or no lifting of it [3]. This type of action has been suggested for uniformly sloped piles of dry homogeneous soil [20], and we expect the optimization process to converge faster with this initialization. The maximum number of iterations is set to 30 with early stopping (patience 3 for tolerance $10^{-4}$). The gradient with respect to $\bm{a}^{\text{load}}$ is calculated by using pytorch.autograd and exploiting that $\psi$ has two input branches, separating $\bm{h}$ and $\bm{a}^{\text{load}}$ so that the encoding part only needs to be evaluated once during the optimization process.
@@ 4.2 V-turn model
A loading cycle involves moving the wheel loader back and forth between the pile and the load receiver in a V-shaped pattern connecting the dump location $\bm{x}^{\text{dump}}$ and dig location $\bm{x}^{\text{dig}}$. We estimate the transportation time and work by numerical integration of an assumed equation of motion for the wheel loader, including its motor strength, variable mass, and rolling resistance. For generating the V-turn paths, we use cubic B-splines [21]. Each V-turn path is formed by the combination of two spline curves. V-turn-1 is composed of the splines connecting $\bm{x}^{\text{dump}}\to\bm{x}^{\text{V1}}$ and $\bm{x}^{\text{V1}}\to\bm{x}^{\text{dig}}$ with condition that the ingoing and outgoing directions should match in the switch back point $\bm{x}^{\text{V1}}$. Similarly, V-turn-2 is composed of two splines connecting $\bm{x}^{\text{dig}}\to\bm{x}^{\text{V2}}$ and $\bm{x}^{\text{V2}}\to\bm{x}^{\text{dump}}$. See Fig. 4 for an illustration.
To compute the splines, we follow the method and notations of [22]. The path to be found is represented by five control points, $\{\bm{p}_{i}\}_{i=0}^{5}$. These are obtained by solving the following equation system
@@
For a given set of endpoints, $\bm{x}^{\text{dump}}$ and $\bm{x}^{\text{dig}}$ (and heading at these points), we compute the connecting splines, via $\bm{x}^{\text{V1}}$ and $\bm{x}^{\text{V2}}$, that minimize the path length and curvature according to the following weighted objective function [22]
@@ 4.4 Look-ahead tree search
To find a near-optimal solution, we develop a look-ahead tree search algorithm that combines the world model and V-turn model for long-horizon predictions. The planning horizon is denoted by integer $N$ and the search depth by $d\leq N$. At each planning step, $n=1,2,\ldots,N$, a finite set of $I$ candidate dig locations $\{\bm{x}_{i}^{\text{dig}}\}_{i=1}^{I}$ and corresponding loading actions $\{\bm{a}_{i}^{\text{load}}\}_{i=1}^{I}$ are considered. We are to select the dig location and loading action that are optimal over the search depth $d$. At the most shallow search depth, $d=1$, this corresponds to picking the candidate that minimizes $\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{n}(\bm{H}_{n},\bm{x}^{\text{dig}}_{n},\bm{a}^{\text{load}}_{n})$. At the first planning step, $n=1$, we know the pile state $\bm{H}_{1}$ and can make this pick. For the future planning steps, we must first expand the pile into its next state from the previous, $\bm{H}_{n+1}=\bm{H}^{\prime}_{n}\equiv\bm{\Phi}(\bm{H}_{n},\bm{x}^{\text{dig}}_{n},\bm{a}^{\text{load}}_{n})$. At larger search depths, $d\geq 2$, we compute the predicted net performance over the search horizon $d$ using the following evaluation function
@@
where $\bar{\bm{x}}_{k}$ and $\bar{\bm{a}}_{k}$ are greedy choices given the pile state $\bar{\bm{H}}_{k}$, that is, those that minimize $\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{k}(\bar{\bm{H}}_{k},\bar{\bm{x}}_{k},\bar{\bm{a}}_{k})$. The pile $\bar{\bm{H}}_{k}$ is, in turn, the result of an expansion from the pile state at the previous search depth, $\bar{\bm{H}}_{k-1}$, using a greedy choice at that level too [^1] If there is a maximum number loading cycles, $N$, then the summation in Eq. (4.4) is limited by this in case $n+d>N$. Also, if $d=1$, there is no second or higher term on the right-hand side.. Evaluating Eq. (4.4) requires that pile state at the top level pile $\bm{H}_{n}$ is expanded into all the candidate piles $\{\bm{H}_{i}^{\prime}\}_{i=1}^{I}$ using $\{\bm{x}_{i}^{\text{dig}}\}_{i=1}^{I}$ and $\{\bm{a}_{i}^{\text{load}}\}_{i=1}^{I}$. The algorithm is displayed in Alg. 1 and pictorially illustrated in Fig. 1.

### v3

# loader_e2e_optimization.bib

@@ Abstract
Wheel loaders in mines and construction sites repeatedly load soil from a pile to load receivers. Automating this task presents a challenging planning problem since each loading’s performance depends on the pile state, which depends on previous loadings. We investigate an end-to-end optimization approach considering future loading outcomes and transportation costs between the pile and load receivers. To predict the evolution of the pile state and the loading performance, we use world models that leverage deep neural networks trained on numerous simulated loading cycles. A look-ahead tree search optimizes the sequence of loading actions by evaluating the performance of thousands of action candidates, which expand into subsequent action candidates under the predicted pile states recursively. Test results demonstrate that, over a horizon of 15 sequential loadings, the look-ahead tree search is 6 % more efficient than a greedy strategy, which always selects the action that maximizes the current single loading performance, and 14 % more efficient than using a fixed loading controller optimized for the nominal case.
@@ 1 Introduction
The task of most construction and mining machines, such as wheel loaders, is to perform long sequences of earthmoving operations in dynamic environments. Optimizing this task involves selecting loading and transportation actions that maximize the gain over time, according to some performance measure. This is challenging for several reasons. First, the performance of a single loading cycle depends highly on the local shape of the pile and the soil properties [singh2006factors]. The loading actions need to be carefully adapted to these conditions. Second, each loading alters the pile shape. From the perspective of future loadings, such a change can be either for the better or worse. There might be actions that are of low performance in the short term but are necessary to achieve higher gains in the future [singh1992task]. Therefore, a greedy strategy of always choosing the loading action that maximizes the performance for a single loading might be sub-optimal in the long run. Such a strategy might gradually deteriorate the pile state or lead to excessive transportation between the dig location and the load receiver. This motivates considering future states when selecting the actions.
We refer to the problem of finding the action sequence that maximizes the total performance of many wheel loading cycles, given the initial pile state, as the wheel loader end-to-end optimization problem. To the best of our knowledge, this problem has not been scientifically explored previously. In fact, until recently, it was computationally intractable. In our previous work [aoshima2023predictor], we developed data-driven models capable of predicting the loading performance and the subsequent state of the pile within a few milliseconds. Model inputs are the previous pile state and selected action parameters. This makes it possible to rapidly evaluate the net performance of numerous alternative action sequences in advance. These models are referred to as a world model, as they predict the new state the environment will transition into, given the previous state and a selected action, and the observations associated with the state transition [ding2024].
@@
We use a look-ahead tree search approach illustrated in Fig. 1. Future action candidates for the foreseen future states are identified recursively, offline. The action candidates are evaluated, using data-driven models developed in previous research [aoshima2023predictor], to compute their respective performance and expand into further action candidates under the new states. The method should apply to any kind of world model with the same prediction capabilities. This approach is, reportedly, similar to how experienced operators intuitively plan their work, but we expect that a computerized model and search algorithms can make more accurate predictions and plan over longer horizons. To understand the importance of the look-ahead search, we examined the effect of using different planning horizons and objectives.
@@
The wheel loader is assumed to have a high-level planner and a low-level control system, which work as the hierarchical architecture described in refs. [lever2001] and [wang2021]. The high-level planner selects a subtask action, such as bucket filling, dumping, or V-shaped transportation, according to the current state of the machine and the work site. A low-level controller performs the subtasks given a motion plan and control parameters. For bucket filling, the machine is equipped with a type of admittance controller [Dobson2017]. Our controller takes four action parameters that determine how reactive the boom and bucket actuation is in response to the momentaneous dig force [aoshima2023predictor]. The performance of a loading cycle is measured in terms of mass, time, and work, with contributions from each subtask.
@@ 2 Related work
The end-to-end optimization problem of sequential loading was discussed in [singh1992task], considered intractable, and substituted with a simpler problem of optimizing each loading cycle by searching in a lower-dimensional action space implicitly constrained by the pile state. An average of 90 % efficiency (bucket payload relative to its capacity) was achieved over 20 cycles. The study was limited to 2D, used a heuristic simulation model for soil displacements and settling, and did not consider time or energy consumption. Attempts were made to improve the performance further by using heuristics to avoid action plans that produce challenging terrain configurations, but the optimization was computationally prohibitive.
@@
What is a good dig location in the pile, given its shape in terms of a heightmap, has been studied from the perspective of both bucket filling and the transportation between the dig and the dump location. In [singh1998multi], a coarse-level planner finds the points along the pile contour closest to the load receiver, and a refined planner was used to avoid curved regions of the pile surface to minimize bucket-side loading and drop-in filling efficiency. A sequential coarse-to-fine planner was studied in [magnusson2015quantitative] and compared to a greedy planner. The coarse planner first produces a sequence of dig regions ordered radially around the pile centroid. The fine planner then selects the dig points inside each region that have the best local pile shape. Simulations (cellular automata) of 50 consecutive loadings showed that the coarse-to-fine strategy maintains a good pile shape and performs at 80–90 % while a greedy strategy (always selecting the dig location with optimal pile shape) initially performs equal or better for ten loadings but drops down to about 60 % after 25 loadings.
@@
Models for adapting the loading control to the pile shape have been studied with different approaches. In [backman2021continuous], reinforcement learning was used to train a multiagent system for autonomous control of an underground load-haul-dump (LHD) vehicle using a depth camera, lidar, and force and kinematics sensors. One agent was used to select a favorable dig position, given the observed pile shape. A second agent had the task of steering the vehicle toward the target position and controlling the forward drive as well as the lift and tilt cylinders to perform bucket filling. The agents were trained in a simulated environment to learn control policies designed for high productivity and energy efficiency over multiple loading cycles while also avoiding collisions and wheel slip. The system achieved bucket filling with 75 % of maximum capacity on average and used 4 % less energy by actively selecting favorable dig locations. The vehicle was confined to operate in a narrow mine drift with little variability in viable dig position or how to navigate to and from the dump location. It is untested how well this approach applies to controlling a wheel loader above ground with much larger action and observation space. The study [lindmark2018computational] explored multiobjective optimization of an LHD control strategy using surrogate models based on data from sequences of simulated loadings in a pile of fragmented rock. Each simulated sequence used a parameterized dig trajectory planner that adapts to the local pile state. Based on many simulations with different parameters, the learned model can identify the loading strategies that ensure high average performance and avoid those that systematically deteriorate the pile state.
@@
Optimization of the V-shaped transportation path between the dig and dump location has been addressed in several studies, focusing on finding the shortest path with curvature consistent with the vehicle’s minimum turning radius [Sarata2005, takei2015simultaneous], taking into account the vehicle dynamics and construction working site constraints [ALSHAER20135315], non-uniform trajectories under forward and reverse driving conditions [SHI2020103570], or minimizing fuel consumption [hong2017path]. Time and energy-efficient tracking of a planned path has been explored in several studies using optimal control [NEZHADALI20161, sardarmehni2023path].
@@
The optimal bucket filling trajectory was researched in [filla2017towards] using the discrete element method (DEM) and used in [frank2018optimal] to find the optimal short loading cycle using dynamic programming. For the sake of computational efficiency, the simulations were limited to quasi-2D piles sloped with the material’s angle of repose. The effect of the pile shape on the optimal trajectory or the resulting 3D pile state has not been explored using DEM simulation, apart from the work in [lindmark2018computational].
@@ 4.1 Loading prediction
We use the data-driven world model developed in [aoshima2023predictor] for a Komatsu WA320-7 wheel loader doing automatic bucket filling in piles of gravel using an admittance controller. The models predict the loading outcome
@@
The models The models are trained and validated on a dataset from more than 10,000 random loading actions on various gravel pile shapes. The trained models, with roughly 10e7 parameters, achieved around 95% accuracy in predicting loading performance and 97% accuracy in predicting the resulting pile state.
As in [aoshima2023predictor], we use different sizes and resolutions of the local heightmap for the pile state and performance predictor models. For the former model, we discretize a quadratic heightmap with 5.2 m sides in a $52\times 52$ grid. For the latter model, we use a $36\times 36$ grid and a side length of 3.6 m.
@@ 4.1.1 Optimal loading action parameters
where $\eta$ is the step length and $\nabla=\partial/\partial\bm{a}^{\text{load}}$. As an initial guess, we use $\bm{a}^{\text{load}}_{n}=[1,0,0.0,0.5,0.5]^{\mathsf{T}}$, which corresponds to the machine thrusting deep into the pile, filling the bucket with little or no lifting of it [aoshima2023predictor]. This type of action has been suggested for uniformly sloped piles of dry homogeneous soil [Bradley1998], and we expect the optimization process to converge faster with this initialization. The maximum number of iterations is set to 30 with early stopping (patience 3 for tolerance $10^{-4}$). The gradient with respect to $\bm{a}^{\text{load}}$ is calculated by using pytorch.autograd and exploiting that $\psi$ has two input branches, separating $\bm{h}$ and $\bm{a}^{\text{load}}$ so that the encoding part only needs to be evaluated once during the optimization process.
@@ 4.2 V-turn model
A loading cycle involves moving the wheel loader back and forth between the pile and the load receiver in a V-shaped pattern connecting the dump location $\bm{x}^{\text{dump}}$ and dig location $\bm{x}^{\text{dig}}$. We estimate the transportation time and work by numerical integration of an assumed equation of motion for the wheel loader, including its motor strength, variable mass, and rolling resistance. For generating the V-turn paths, we use cubic B-splines [piegl1996nurbs]. In narrow spaces, Dubins curves may be the preferred choice to ensure that the vehicle’s minimum turning radius is respected [dubins1957curves]. Each V-turn path is formed by the combination of two spline curves. V-turn-1 is composed of the splines connecting $\bm{x}^{\text{dump}}\to\bm{x}^{\text{V1}}$ and $\bm{x}^{\text{V1}}\to\bm{x}^{\text{dig}}$ with condition that the ingoing and outgoing directions should match in the switch back point $\bm{x}^{\text{V1}}$. Similarly, V-turn-2 is composed of two splines connecting $\bm{x}^{\text{dig}}\to\bm{x}^{\text{V2}}$ and $\bm{x}^{\text{V2}}\to\bm{x}^{\text{dump}}$. See Fig. 4 for an illustration.
To compute the splines, we follow the method and notations of [MAEKAWA2010350]. The path to be found is represented by five control points, $\{\bm{p}_{i}\}_{i=0}^{5}$. These are obtained by solving the following equation system
@@
For a given set of endpoints, $\bm{x}^{\text{dump}}$ and $\bm{x}^{\text{dig}}$ (and heading at these points), we compute the connecting splines, via $\bm{x}^{\text{V1}}$ and $\bm{x}^{\text{V2}}$, that minimize the path length and curvature according to the following weighted objective function [MAEKAWA2010350]
@@ 4.4 Look-ahead tree search
To find a near-optimal solution, we develop a look-ahead tree search algorithm that combines the world model and V-turn model for long-horizon predictions. The planning horizon is denoted by integer $N$ and the search depth by $d\leq N$. At each planning step, $n=1,2,\ldots,N$, a finite set of $I$ candidate dig locations $\{\bm{x}_{i}^{\text{dig}}\}_{i=1}^{I}$ and corresponding loading actions $\{\bm{a}_{i}^{\text{load}}\}_{i=1}^{I}$ are considered. We are to select the dig location and loading action that are optimal over the search depth $d$. At the most shallow search depth, $d=1$, this corresponds to picking the candidate that minimizes $\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{n}(\bm{H}_{n},\bm{x}^{\text{dig}}_{n},\bm{a}^{\text{load}}_{n})$. At the first planning step, $n=1$, we know the pile state $\bm{H}_{1}$ and can make this pick. For the future planning steps, we must first expand the pile into its next state from the previous, $\bm{H}_{n+1}=\bm{H}^{\prime}_{n}\equiv\bm{\Phi}(\bm{H}_{n},\bm{x}^{\text{dig}}_{n},\bm{a}^{\text{load}}_{n})$. At larger search depths, $d\geq 2$, we compute the predicted net performance over the search horizon $d$ using the following approximation of the evaluation function
@@
where $\bar{\bm{x}}_{k}$ and $\bar{\bm{a}}_{k}$ are greedy choices given the pile state $\bar{\bm{H}}_{k}$, that is, those that minimize $\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{k}(\bar{\bm{H}}_{k},\bar{\bm{x}}_{k},\bar{\bm{a}}_{k})$. The pile $\bar{\bm{H}}_{k}$ is, in turn, the result of an expansion from the pile state at the previous search depth, $\bar{\bm{H}}_{k-1}$, using a greedy choice at that level too [^1] If there is a maximum number loading cycles, $N$, then the summation in Eq. (4.4) is limited by this in case $n+d>N$. Also, if $d=1$, there is no second or higher term on the right-hand side.. This approximation is to avoid doing an exhaustive search over all action candidates. Evaluating Eq. (4.4) requires that pile state at the top level pile $\bm{H}_{n}$ is expanded into all the candidate piles $\{\bm{H}_{i}^{\prime}\}_{i=1}^{I}$ using $\{\bm{x}_{i}^{\text{dig}}\}_{i=1}^{I}$ and $\{\bm{a}_{i}^{\text{load}}\}_{i=1}^{I}$. The algorithm is displayed in Alg. 1 and pictorially illustrated in Fig. 1.